How many liters of ammonia (NH3) are contained in 800 ml of a solution that is 12% ammonia by volume?


1
Expert's answer
2022-05-11T10:40:04-0400

Solution:

Percent by volume = (Volume of the solute / Volume of the solution) × 100%

solute = NH3

Therefore,

12% = (Volume of NH3 / 0.8 L) × 100%

Volume of NH3 / 0.8 L = 0.12

Volume of NH3 = 0.12 × 0.8 L = 0.096 L

Volume of NH3 = 0.096 L


Answer: 0.096 liters of ammonia (NH3)